A (n)=0.5*3^n-1. I need to find the first fourth and eighth term. If you explain in simplest terms I would be super greatful
Found 2 solutions by solver91311, Boreal:
Answer by solver91311(24713)



Write in pencil.

Erase the . Write in its place. Do the indicated arithmetic.

Repeat the process for and . Hint: Anything to the zero power is 1.

John

My calculator said it, I believe it, that settles it

Answer by Boreal(15235)

A (n)=0.5*3^(n-1); I am assuming n-1 is the exponent.
I would substitute the number for n
A(1)=0.5*3^(1-1)=0.5*3^0; 3^0=1, so A(1)=0.5
A(4)=0.5*3^(4-1)=0.5*3^3; 3^3=27, and 27*(0.5)=13.5
A(8)=0.5*3^(8-1)=0.5*3^7;3^7=2187. and 2187*0.5=1093.5